# Domino tilings, nonintersecting lattice paths and subclasses of Koutschan–Krattenthaler–Schlosser determinants

### v1

@@ 1.1. Background
It is notable that the convention used in the original paper of Di Francesco [6] was slightly different; he also put $\binom{\alpha}{p}=0$ for $-1\leq\alpha<p$. From (1.2), Di Francesco further conjectured a product formula [6, p. 43, Conjecture 8.1] for $\cyD(n)$:
@@
To connect Di Francesco’s binomial determinant (1.2) with the KKS determinants given in (1.1), we first recall the work of Corteel, Huang and Krattenthaler [4] on domino tilings of generalized Aztec triangles, in which Di Francesco’s setting is extended. In particular, it was shown in [4, Theorem 4.1, eq. (4.1)] that for nonnegative integers $k$, the number (${}_{k}(n)$ of domino tilings of a certain generalized Aztec triangle (see Remark 2.4 for a concrete description) admits the product form
(${}_{k}(n)=\prod_{j=1}^{k-1}(2j+1)^{j-k}\prod_{i\geq 0}\left(\prod_{s=-2k+4i+1}^{-k+2i}(2n+s)\prod_{s=k-2i}^{2k-4i-2}(2n+s)\right),$ (1.4)
@@
(${}_{n}(n+\delta)=\frac{1}{2}\cdot\underset{{0\leq i,j\leq n-1}}{\det}\left(2^{i}\binom{2j+i+2\delta+1}{2j+1}+\binom{2j-i+2\delta+1}{2j+1}\right),$ (1.5)
@@
where $\Gamma(x)$ is the gamma function. Therefore, the counting function (${}_{n}(n+\delta)$ for domino tilings forms a subclass of KKS determinants in a combinatorial context; this is especially adorable because of the melody played by the above product-like evaluation.
Noting also that after certain modifications, some of their product formulas ([16, p. 12, Theorem 6, eq. (5.3)] and [16, p. 22, Theorem 15, eq. (8.9)]) appear coincidentally in the context of $(-1)$-enumerations of arrowed Gelfand–Tsetlin patterns introduced by Fischer and Schreier-Aigner ([5, p. 3, Theorem 1] and [5, p. 4, Theorem 2]), Koutschan, Krattenthaler and Schlosser proposed a request at the end of their paper [16, p. 32] for combinatorial interpretations of their determinants. Here we remark that the connection between KKS determinants and Fischer and Schreier-Aigner’s arrowed Gelfand–Tsetlin patterns is still unclear.
@@ 1.2. Main results
Briefly speaking, we are going to count domino tilings of two types of domains generalized by Di Francesco’s Aztec triangles [6]. In particular, such domains will be characterized by an arithmetic progression $(s(n-1)+r,s(n-2)+r,\ldots,r)$ for certain $s,r\in\mathbb{N}$. Furthermore, we assign three weights $w_{1},w_{2},w_{3}$ to different types of dominoes. Such weighted enumerations of domino tilings of the two types of domains are denoted by (${}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ and ^$\widehat{\cyD}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$, respectively; see (2.8) and (2.9) for their concrete definitions.
Meanwhile, in light of a bijection due to Corteel, Huang and Krattenthaler [4], we will also consider nonintersecting Delannoy and H-Delannoy paths, respectively, while these nonintersecting paths will be characterized by the same arithmetic progression with parameters $s,r\in\mathbb{N}$. Once again, we assign weights $w_{1},w_{2},w_{3}$ to three types of steps in the paths and conduct similar weighted enumerations by the counting functions (${}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ and ^$\widehat{\cyL}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ defined in (2.3) and (2.4), respectively.
@@
(${}_{a,0,a,a}^{(m,l)}(n)=\underset{{0\leq i,j\leq n-1}}{\det}\left(l^{j}\binom{mi+j+a}{mi+a}+\binom{mi-j+a}{mi+a}\right).$
@@
(${}_{a+1,1,a+1,a-1}^{(m,l)}(n)=\underset{{0\leq i,j\leq n-1}}{\det}\left(l^{j+1}\binom{mi+j+a+1}{mi+a+1}+\binom{mi-j+a-1}{mi+a+1}\right).$
@@
(${}_{1,1,1}^{(1,0)}(n)$ $=\tfrac{1}{2}{}_{0,0,0,0}^{(2,2)}(n)=2^{n(n-1)}\prod_{i=1}^{n}\frac{\Gamma(i)\Gamma(2i)\Gamma(4i-3)\Gamma(\frac{3i-1}{2})}{\Gamma(2i-1)\Gamma(3i-1)\Gamma(3i-2)\Gamma(\frac{i+1}{2})},$ (1.12)
(${}_{1,1,1}^{(1,1)}(n)$ $=\tfrac{1}{2}{}_{1,0,1,1}^{(2,2)}(n)=2^{n(n-1)}\prod_{i=1}^{n}\frac{\Gamma(i)\Gamma(4i-1)\Gamma(\frac{3i-2}{2})}{\Gamma(3i)\Gamma(3i-2)\Gamma(\frac{i}{2})},$ (1.13)
(${}_{1,3,1}^{(1,1)}(n)$ $=\tfrac{1}{2}{}_{1,0,1,1}^{(2,4)}(n)=3^{\frac{1}{2}n(n-1)}\prod_{i=1}^{n}\frac{\Gamma(3i-1)\Gamma(\frac{i+1}{2})}{\Gamma(2i)\Gamma(\frac{3i-1}{2})},$ (1.14)
(${}_{1,2,1}^{(2,0)}(n)$ $=\tfrac{1}{2}{}_{0,0,0,0}^{(3,3)}(n)=2^{\frac{1}{2}n(n-1)}\prod_{i=1}^{n}\frac{\Gamma(4i-3)\Gamma(\frac{i+1}{3})}{\Gamma(3i-2)\Gamma(\frac{4i-2}{3})},$ (1.15)
(${}_{1,2,1}^{(2,1)}(n)$ $=\tfrac{1}{2}{}_{1,0,1,1}^{(3,3)}(n)=2^{\frac{1}{2}n(n-3)}3^{-n}\prod_{i=1}^{n}\frac{\Gamma(4i-1)\Gamma(\frac{i}{3})}{\Gamma(3i-1)\Gamma(\frac{4i}{3})},$ (1.16)
(${}_{1,2,1}^{(2,2)}(n)$ $=\tfrac{1}{2}{}_{2,0,2,2}^{(3,3)}(n)=2^{\frac{1}{2}n(n-5)}\prod_{i=1}^{n}\frac{\Gamma(4i+1)\Gamma(\frac{i+2}{3})}{\Gamma(3i+1)\Gamma(\frac{4i+2}{3})}.$ (1.17)
@@
^$\widehat{\cyD}_{1,1,1}^{(1,0)}(n)$ =($={}_{1,1,1,-1}^{(2,2)}(n)=2^{n(n-1)}\prod_{i=1}^{n}\frac{\Gamma(i)\Gamma(4i-1)\Gamma(\frac{3i-2}{2})}{\Gamma(3i)\Gamma(3i-2)\Gamma(\frac{i}{2})},$ (1.18)
^$\widehat{\cyD}_{1,1,1}^{(1,1)}(n)$ =($={}_{2,1,2,0}^{(2,2)}(n)=\prod_{i=1}^{n}\frac{\Gamma(4i)\Gamma(\frac{i+2}{2})}{\Gamma(3i)\Gamma(\frac{3i+2}{2})},$ (1.19)
^$\widehat{\cyD}_{1,3,1}^{(1,0)}(n)$ =($={}_{1,1,1,-1}^{(2,4)}(n)=3^{\frac{1}{2}n(n+1)}\prod_{i=1}^{n}\frac{\Gamma(3i-1)\Gamma(\frac{i+1}{2})}{\Gamma(2i)\Gamma(\frac{3i-1}{2})}.$ (1.20)
@@
(${}_{1,1,1}^{(3,3)}(n)=\tfrac{1}{2}{}_{3,0,3,3}^{(4,2)}(n)=2\prod_{i=1}^{n}\frac{\Gamma(6i-1)\Gamma(\frac{i+3}{4})}{\Gamma(5i)\Gamma(\frac{5i-1}{4})},$ (1.21)
@@
^$\widehat{\cyD}_{1,1,1}^{(3,1)}(n)={}_{2,1,2,0}^{(4,2)}(n)=\frac{\Gamma(n+1)}{\Gamma(2n+1)}\prod_{i=1}^{n}\frac{\Gamma(6i-1)\Gamma(\frac{i+2}{4})}{\Gamma(5i-1)\Gamma(\frac{5i-2}{4})}.$ (1.22)
@@ 2.1. Nonintersecting lattice paths
A system of nonintersecting Delannoy paths marked by $\lambda$ is an $n$-tuple $p=(p_{1},\ldots,p_{n})$ of Delannoy paths such that $p_{j}$ moves from $(-j,j)$ to $(\lambda_{j}-j,n)$ for $1\leq j\leq n$ and that these Delannoy paths are pairwise nonintersecting in the sense that every two different paths do not have any lattice point in common. We denote by $\mydutchcal{L}_{\lambda}(n)$ the set of such nonintersecting Delannoy paths.
Similarly, a system of nonintersecting H-Delannoy paths marked by $\lambda$ is an $n$-tuple $\hat{p}=(\hat{p}_{1},\ldots,\hat{p}_{n})$ of H-Delannoy paths such that $\hat{p}_{j}$ moves from $(-j,j)$ to $(\lambda_{j}-j,n+1)$ for $1\leq j\leq n$ and that these H-Delannoy paths are pairwise nonintersecting. The set of nonintersecting H-Delannoy paths marked by $\lambda$ is written as $\widehat{\mydutchcal{L}}_{\lambda}(n)$.
@@
(${}_{w_{1},w_{2},w_{3}}^{\lambda}(n)$ $:=\sum_{p\in\mydutchcal{L}_{\lambda}(n)}w_{\mathrm{S}}(p),$
^$\widehat{\cyL}_{w_{1},w_{2},w_{3}}^{\lambda}(n)$ $:=\sum_{\hat{p}\in\widehat{\mydutchcal{L}}_{\lambda}(n)}w_{\mathrm{S}}(\hat{p}).$
@@
#$\#_{p}(\rightarrow)+\#_{p}(\nearrow)=\lambda_{1}+\lambda_{2}+\cdots+\lambda_{n}=|\lambda|,$
@@
#$\#_{p}(\uparrow)+\#_{p}(\nearrow)=(n-1)+(n-2)+\cdots+(n-n)=\tbinom{n}{2}.$
@@
#$\#_{\hat{p}}(\rightarrow)+\#_{\hat{p}}(\nearrow)=\lambda_{1}+\lambda_{2}+\cdots+\lambda_{n}=|\lambda|,$
@@
#$\#_{\hat{p}}(\uparrow)+\#_{\hat{p}}(\nearrow)=(n-0)+(n-1)+\cdots+(n-(n-1))=\tbinom{n+1}{2}.$
@@
(${}_{cw_{1},w_{2},cw_{3}}^{\lambda}(n)$ $=c^{|\lambda|}{}_{w_{1},w_{2},w_{3}}^{\lambda}(n),$
(${}_{w_{1},cw_{2},cw_{3}}^{\lambda}(n)$ $=c^{\binom{n}{2}}{}_{w_{1},w_{2},w_{3}}^{\lambda}(n).$
@@
^$\widehat{\cyL}_{cw_{1},w_{2},cw_{3}}^{\lambda}(n)$ $=c^{|\lambda|}\widehat{\cyL}_{w_{1},w_{2},w_{3}}^{\lambda}(n),$
^$\widehat{\cyL}_{w_{1},cw_{2},cw_{3}}^{\lambda}(n)$ $=c^{\binom{n+1}{2}}\widehat{\cyL}_{w_{1},w_{2},w_{3}}^{\lambda}(n).$
@@
Let $\0c_{1},c_{2}\in\mathbb{C}\backslash\{0\}$. Then in view of Proposition 2.1, we have
(${}_{c_{1},(l-1)c_{2},c_{1}c_{2}}^{(m-1,a)}(n)$ $=c_{1}^{(m-1)\binom{n}{2}+an}c_{2}^{\binom{n}{2}}{}_{1,l-1,1}^{(m-1,a)}(n),$ (2.5)
^$\widehat{\cyL}_{c_{1},(l-1)c_{2},c_{1}c_{2}}^{(m-1,a)}(n)$ $=c_{1}^{(m-1)\binom{n}{2}+an}c_{2}^{\binom{n}{2}+n}\widehat{\cyL}_{1,l-1,1}^{(m-1,a)}(n).$ (2.6)
So the KKS determinants in (1.10) and (1.11), up to a scalar, may also be related to our enumerations of nonintersecting lattice paths counted by other choices of weights as indicated on the left of (2.5) and (2.6). A particularly interesting case here is to choose $c_{1}=l-1$ and $c_{2}=1/(l-1)$ when $l\neq 1$, so that we arrive at expressions in terms of KKS determinants for (${}_{l-1,1,1}^{(m-1,a)}(n)$ and ^$\widehat{\cyL}_{l-1,1,1}^{(m-1,a)}(n)$.

### v2

@@ 1.1. Background
It is notable that the convention used in the original paper of Di Francesco [6] was slightly different; he also put $\binom{\alpha}{p}=0$ for $1\leq\alpha<p$. From (1.2), Di Francesco further conjectured a product formula [6, p. 43, Conjecture 8.1] for $\cyD(n)$:
@@
[…]
${}_{k}(n)=\prod_{j=1}^{k-1}(2j+1)^{j-k}\prod_{i\geq 0}\left(\prod_{s=-2k+4i+1}^{-k+2i}(2n+s)\prod_{s=k-2i}^{2k-4i-2}(2n+s)\right),$ (1.4)
@@
${}_{n}(n+\delta)=\frac{1}{2}\cdot\underset{{0\leq i,j\leq n-1}}{\det}\left(2^{i}\binom{2j+i+2\delta+1}{2j+1}+\binom{2j-i+2\delta+1}{2j+1}\right)$, (1.5)
@@
[…]
Noting also that after certain modifications, some of their product formulas ([16, p. 12, Theorem 6, eq. (5.3)] and [16, p. 22, Theorem 15, eq. (8.9)]) appear coincidentally in the context of (1)(-1)-enumerations of arrowed Gelfand–Tsetlin patterns introduced by Fischer and Schreier-Aigner ([5, p. 3, Theorem 1] and [5, p. 4, Theorem 2]), Koutschan, Krattenthaler and Schlosser proposed a request at the end of their paper [16, p. 32] for combinatorial interpretations of their determinants. Here we remark that the connection between KKS determinants and Fischer and Schreier-Aigner’s arrowed Gelfand–Tsetlin patterns is still unclear.
@@ 1.2. Main results
Briefly speaking, we are going to count domino tilings of two types of domains generalized by Di Francesco’s Aztec triangles [6]. In particular, such domains will be characterized by an arithmetic progression $(s(n-1)+r,s(n-2)+r,\ldots,r)$ for certain $s,r\in\mathbb{N}$. Furthermore, we assign three weights $w_{1},w_{2},w_{3}$ to different types of dominoes. Such weighted enumerations of domino tilings of the two types of domains are denoted by $_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ and $\widehat{\cyD}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$, respectively; see (2.8) and (2.9) for their concrete definitions.
Meanwhile, in light of a bijection due to Corteel, Huang and Krattenthaler [4], we will also consider nonintersecting Delannoy and H-Delannoy paths, respectively, while these nonintersecting paths will be characterized by the same arithmetic progression with parameters $s,r\in\mathbb{N}$. Once again, we assign weights $w_{1},w_{2},w_{3}$ to three types of steps in the paths and conduct similar weighted enumerations by the counting functions $_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ and $\widehat{\cyL}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ defined in (2.3) and (2.4), respectively.
@@
${}_{a,0,a,a}^{(m,l)}(n)=\underset{{0\leq i,j\leq n-1}}{\det}\left(l^{j}\binom{mi+j+a}{mi+a}+\binom{mi-j+a}{mi+a}\right)$.
@@
${}_{a+1,1,a+1,a-1}^{(m,l)}(n)=\underset{{0\leq i,j\leq n-1}}{\det}\left(l^{j+1}\binom{mi+j+a+1}{mi+a+1}+\binom{mi-j+a-1}{mi+a+1}\right)$.
@@
${}_{1,1,1}^{(1,0)}(n)$ $=\tfrac{1}{2}{}_{0,0,0,0}^{(2,2)}(n)=2^{n(n-1)}\prod_{i=1}^{n}\frac{\Gamma(i)\Gamma(2i)\Gamma(4i-3)\Gamma(\frac{3i-1}{2})}{\Gamma(2i-1)\Gamma(3i-1)\Gamma(3i-2)\Gamma(\frac{i+1}{2})},$ (1.12)
${}_{1,1,1}^{(1,1)}(n)$ $=\tfrac{1}{2}{}_{1,0,1,1}^{(2,2)}(n)=2^{n(n-1)}\prod_{i=1}^{n}\frac{\Gamma(i)\Gamma(4i-1)\Gamma(\frac{3i-2}{2})}{\Gamma(3i)\Gamma(3i-2)\Gamma(\frac{i}{2})},$ (1.13)
${}_{1,3,1}^{(1,1)}(n)$ $=\tfrac{1}{2}{}_{1,0,1,1}^{(2,4)}(n)=3^{\frac{1}{2}n(n-1)}\prod_{i=1}^{n}\frac{\Gamma(3i-1)\Gamma(\frac{i+1}{2})}{\Gamma(2i)\Gamma(\frac{3i-1}{2})},$ (1.14)
${}_{1,2,1}^{(2,0)}(n)$ $=\tfrac{1}{2}{}_{0,0,0,0}^{(3,3)}(n)=2^{\frac{1}{2}n(n-1)}\prod_{i=1}^{n}\frac{\Gamma(4i-3)\Gamma(\frac{i+1}{3})}{\Gamma(3i-2)\Gamma(\frac{4i-2}{3})},$ (1.15)
${}_{1,2,1}^{(2,1)}(n)$ $=\tfrac{1}{2}{}_{1,0,1,1}^{(3,3)}(n)=2^{\frac{1}{2}n(n-3)}3^{-n}\prod_{i=1}^{n}\frac{\Gamma(4i-1)\Gamma(\frac{i}{3})}{\Gamma(3i-1)\Gamma(\frac{4i}{3})},$ (1.16)
${}_{1,2,1}^{(2,2)}(n)$ $=\tfrac{1}{2}{}_{2,0,2,2}^{(3,3)}(n)=2^{\frac{1}{2}n(n-5)}\prod_{i=1}^{n}\frac{\Gamma(4i+1)\Gamma(\frac{i+2}{3})}{\Gamma(3i+1)\Gamma(\frac{4i+2}{3})}.$ (1.17)
@@
$\widehat{\cyD}_{1,1,1}^{(1,0)}(n)$ $={}_{1,1,1,-1}^{(2,2)}(n)=2^{n(n-1)}\prod_{i=1}^{n}\frac{\Gamma(i)\Gamma(4i-1)\Gamma(\frac{3i-2}{2})}{\Gamma(3i)\Gamma(3i-2)\Gamma(\frac{i}{2})},$ (1.18)
$\widehat{\cyD}_{1,1,1}^{(1,1)}(n)$ $={}_{2,1,2,0}^{(2,2)}(n)=\prod_{i=1}^{n}\frac{\Gamma(4i)\Gamma(\frac{i+2}{2})}{\Gamma(3i)\Gamma(\frac{3i+2}{2})},$ (1.19)
$\widehat{\cyD}_{1,3,1}^{(1,0)}(n)$ $={}_{1,1,1,-1}^{(2,4)}(n)=3^{\frac{1}{2}n(n+1)}\prod_{i=1}^{n}\frac{\Gamma(3i-1)\Gamma(\frac{i+1}{2})}{\Gamma(2i)\Gamma(\frac{3i-1}{2})}.$ (1.20)
@@
${}_{1,1,1}^{(3,3)}(n)=\tfrac{1}{2}{}_{3,0,3,3}^{(4,2)}(n)=\prod_{i=1}^{n}\frac{\Gamma(6i-1)\Gamma(\frac{i+3}{4})}{\Gamma(5i)\Gamma(\frac{5i-1}{4})},$ (1.21)
@@
$\widehat{\cyD}_{1,1,1}^{(3,1)}(n)={}_{2,1,2,0}^{(4,2)}(n)=\frac{\Gamma(n+1)}{\Gamma(2n+1)}\prod_{i=1}^{n}\frac{\Gamma(6i-1)\Gamma(\frac{i+2}{4})}{\Gamma(5i-1)\Gamma(\frac{5i-2}{4})}.$ (1.22)
@@ 2.1. Nonintersecting lattice paths
A system of nonintersecting Delannoy paths marked by $\lambda$ is an $n$-tuple $p=(p_{1},\ldots,p_{n})$ of Delannoy paths such that $p_{j}$ moves from (j,j)(-j,j) to $(\lambda_{j}-j,n)$ for $1\leq j\leq n$ and that these Delannoy paths are pairwise nonintersecting in the sense that every two different paths do not have any lattice point in common. We denote by $\mydutchcal{L}_{\lambda}(n)$ the set of such nonintersecting Delannoy paths.
Similarly, a system of nonintersecting H-Delannoy paths marked by $\lambda$ is an $n$-tuple $\hat{p}=(\hat{p}_{1},\ldots,\hat{p}_{n})$ of H-Delannoy paths such that $\hat{p}_{j}$ moves from (j,j)(-j,j) to $(\lambda_{j}-j,n+1)$ for $1\leq j\leq n$ and that these H-Delannoy paths are pairwise nonintersecting. The set of nonintersecting H-Delannoy paths marked by $\lambda$ is written as $\widehat{\mydutchcal{L}}_{\lambda}(n)$.
@@
[…]
@@
$\#_{p}(\rightarrow)+\#_{p}(\nearrow)=\lambda_{1}+\lambda_{2}+\cdots+\lambda_{n}=|\lambda|,$
@@
$\#_{p}(\uparrow)+\#_{p}(\nearrow)=(n-1)+(n-2)+\cdots+(n-n)=\tbinom{n}{2}$.
@@
$\#_{\hat{p}}(\rightarrow)+\#_{\hat{p}}(\nearrow)=\lambda_{1}+\lambda_{2}+\cdots+\lambda_{n}=|\lambda|,$
@@
$\#_{\hat{p}}(\uparrow)+\#_{\hat{p}}(\nearrow)=(n-0)+(n-1)+\cdots+(n-(n-1))=\tbinom{n+1}{2}$.
@@
[…]
@@
$\widehat{\cyL}_{cw_{1},w_{2},cw_{3}}^{\lambda}(n)$ $=c^{|\lambda|}\widehat{\cyL}_{w_{1},w_{2},w_{3}}^{\lambda}(n),$
$\widehat{\cyL}_{w_{1},cw_{2},cw_{3}}^{\lambda}(n)$ $=c^{\binom{n+1}{2}}\widehat{\cyL}_{w_{1},w_{2},w_{3}}^{\lambda}(n).$
@@
[…]
${}_{c_{1},(l-1)c_{2},c_{1}c_{2}}^{(m-1,a)}(n)$ $=c_{1}^{(m-1)\binom{n}{2}+an}c_{2}^{\binom{n}{2}}{}_{1,l-1,1}^{(m-1,a)}(n),$ (2.5)
$\widehat{\cyL}_{c_{1},(l-1)c_{2},c_{1}c_{2}}^{(m-1,a)}(n)$ $=c_{1}^{(m-1)\binom{n}{2}+an}c_{2}^{\binom{n}{2}+n}\widehat{\cyL}_{1,l-1,1}^{(m-1,a)}(n).$ (2.6)
So the KKS determinants in (1.10) and (1.11), up to a scalar, may also be related to our enumerations of nonintersecting lattice paths counted by other choices of weights as indicated on the left of (2.5) and (2.6). A particularly interesting case here is to choose $c_{1}=l-1$ and $c_{2}=1/(l-1)$ when $l\neq 1$, so that we arrive at expressions in terms of KKS determinants for $_{l-1,1,1}^{(m-1,a)}(n)$ and $\widehat{\cyL}_{l-1,1,1}^{(m-1,a)}(n)$.